OpenStudy (anonymous):

The price of a certain monthly magazine is x dollars per copy. The total profit from the sales of the magazine is P dollars. It is given that P=Y+Z, where Y varies directly as x and Z varies directly as the square of x. When x is 20, P is 80000; when x is 35, P is 87500. (a) Express P in terms of x. ==>P=6000x-100x^(2) (b) Find the profit when x=15. ==>67500 (c) Using the method of completing the square, express P in the form a-b(x-c)^(2), where a,b and c are constants. Find the values of a,b and c. ==>??? (d) Hence, or otherwise, find the value of x when P is maximum. ==>???

OpenStudy (phi):

(a) and (b) are good. for part (c) you start with \[ -100 x^2 +6000 x = P \] for how to complete the square see http://www.khanacademy.org/math/algebra/quadratics/completing_the_square/v/completing-the-square

OpenStudy (phi):

but in the meantime, we want the x^2 term to have a coefficient of 1 (because this makes "completing the square" easier) one way is divide both sides of the equation (all terms) by -100 \[ x^2 -60x = \frac{P}{-100} \] the next step is divide the number in front of the x by 2: -60/2= -30 (see the video for why we do this) remember that -30 next square the -30 to get 900 we add this 900 to both sides of the equation \[ x^2 -60x +900 = \frac{P}{-100} +900\] now it may not be obvious, but the stuff on the left side of the = is the same as (x-30)^2 (that is why we want to remember the -30) As a check, use FOIL: (x-30)*(x-30)= x^2 -30x -30x +900= x^2-60x+900 so we now have \[ (x-30)^2 = \frac{P}{-100} +900\]

OpenStudy (phi):

They want express P in the form a-b(x-c)^(2), where a,b and c are constants. Find the values of a,b and c. \[ (x-30)^2 = \frac{P}{-100} +900 \] so let's solve for P. add -900 to both sides \[ \frac{P}{-100} +900 -900 = -900 + (x-30)^2 \] or \[ \frac{P}{-100} = -900 + (x-30)^2 \] multiply both sides by -100: \[ P= 90000 - 100 (x-30)^2 \] that is the from they want, where a is 90,000 b is -100 and c is 30

OpenStudy (phi):

**b is +100 because they use the form a-b(x-c)^(2) and we have 90000 - 100(x -30)^2

OpenStudy (phi):

to do (d), look at the equation P= 90000 - 100 (x-30)^2 notice it is 90,000 MINUS something. for P to be as large as possible we want to subtract as little as possible. the term 100*(x-30)^2 is either 0 or positive (because even if x-30 is a minus, after squaring it, it will be positive. The smallest it can be is zero. when it is zero, P will be the largest it can be. what x value makes 100*(x-30)^2 zero ?